Prove that a sequence is bounded if and only if it is both bounded above and bounded below


1
Expert's answer
2021-02-04T06:32:12-0500

Let "<"an">" be a sequence. given that "<"an">" is bounded, then by definition of boundedness

"\exist" k "\in" R such that |an| "\leq" k "\forall" n

"\implies" -k "\leq" |an| "\leq" k "\forall" n

"\implies" -k "\leq" |an| & |an| "\leq" k "\forall" n

which implies that "<"an">" is bounded below by -k and bounded above by k.

therefore, "<"an">" is both bounded above and bounded below.

Conversely:

now suppose that "<"an">" is both bounded above and bounded below.

let L, k "\isin" R and "<"an">" is bounded below by L and bounded above by k for every n

"\implies" an "\leq" k & an"\geq" L "\forall" n

"\because" -|L| - |k| "\leq" L & k"\leq" |L| + |k|

then,

if L "\leq" an "\leq" k "\forall" n

"\implies" -|L| - |k| "\leq" L "\leq" |an| "\leq" k "\leq" |L| + |k| "\forall" n

"\implies" -|L| - |k| "\leq" |an| "\leq" |L| + |k| "\forall" n

hence we conclude that "<"an">" is bounded by |L| + |k|.

"<"an">" is bounded.